OpenStudy (thatonegirl_):

If f is a differentiable function, then f'(a) is given by which of the following without further restrictions on f?

OpenStudy (thatonegirl_):

\[\lim_{h \rightarrow 0}\frac{ f(a+h)-f(a) }{ h}\] \[\lim_{x \rightarrow a} \frac{ f(x)-f(a) }{ x-a}\] \[\lim_{x \rightarrow a}\frac{ f(x+h)-f(x) }{ h }\]

OpenStudy (thatonegirl_):

Oh and the options are A only B only A,B, and C A and B only A and C only